OpenStudy (anonymous):

write a polynomial function in standard form with the coefficent whose zero include 4, 5i and -5i is f(x)=x^3

OpenStudy (anonymous):

First you have to find your zeros. Your zeros are the opposite of 4,5i, and -5i then put it in an equation. it should be like this: \[(x-4)(x-1+5i)(x-1-5i)\] then you just times them all together. You should multiply the imaginary numbers first so they cancel.(keep in mind i^2=-1 so 5i(-5i)=-25(i^2)=25 you should get: \[x^2-2x+26\] then multiply that with x-4. you should get this:\[x^3-6x^2+34x-104\]